# Exploring and Learning Structure: Active Inference Approach in Navigational Agents

### v1

@@ 1 Introduction
Building on these concepts, we introduce a novel model that dynamically learns environmental structure and expands its cognitive map. Integrating visual information and proprioception, our model constructs locations and connections within its cognitive map. Starting with uncertainty, the model envisions action outcomes, expanding its map by incorporating hypotheses into its generative model, analogous to Bayesian model reduction [8] that grows its model upon receiving new observation, we extend ours upon predicted beliefs. This process allows our model to efficiently navigate and comprehend environmental structures with minimal steps, using an active inference navigation scheme [16]. Compared to the Clone-Structured Graph (CSCG) [23], our model rapidly learns environmental layouts more efficiently. An exploration run is shown in Figure 1, displaying from left to right the full environment, the extracted observation and exploration path and the resulting internal map of the agent.
@@ 3.1 Inference and spatial abstraction
The classical inference scheme heavily relies on past and current experiences to localise the agent within its environment, using observation alone, the agent would be weak to aliased observations at different locations. By combining observation with the agent’s proprioception the model is much more robust in differentiating ambiguous environments. The internal positioning $p_{0}$ is initialised at the start of exploration in the absence of prior information, and is updated as the agent transitions between rooms (e.i., by passing through a door), as illustrated in Figure 2, thus as long as the agent is confident in its current state.
@@
Figure 2: POMDP of our generative model transitioning from past and present (up to time-step $t$) to future (from time-step $t+1$). The state $s_{t}$ and pose $p_{t}$ are determined by the corresponding observation $o_{t}$ and influenced by the previous state $s_{t-1}$, pose $p_{t-1}$ and action $a_{t-1}$. Past actions and observations are assumed observable, indicated by a blue colour. In the future, the actions are defined by a policy $\pi$ influencing the new states and position in orange and new predictions in grey.
@@ 3.3 Incorporating spatial dynamics in model parameters
To extend the internal map (our state space), we propose a novel approach where the agent predicts one-step policy outcomes in all directions. $B^{\prime}$ can expand its position dimension given a motion and A’ considers the probability of being at a given state given the position. This results in the following alternative joint probability :
$A^{\prime}*B^{\prime}=P(s_{t+1}|p_{t+1})P(p_{t+1}|p_{t},a_{t})$ (4)
As can be foreseen, if the expected motion leads to an un-visited location, $s_{t+1}$ does not exist in the model. The state is undefined while the position $p_{t+1}$ is certain, therefore all Markov matrices are expected to grow in their state dimension to match this new prediction. This process enables the dynamic expansion of the dimensions of both the observation and position likelihoods ($A$, $A^{\prime}$) and state transition ($B$) to consider the novel state $s_{t+1}$ in a process equivalent to [8]. Subsequently, the state transition probability ($B$) and position likelihood ($A^{\prime}$) can be updated with the imagined beliefs, forming new connections enriching the agent’s ability to reason about potential unexplored areas within the environment, assisting in decision-making processes. While $s_{t+1}$ can be predicted, it lacks information regarding the specific observation $o_{t+1}$ expected in that location. Such areas exhibit high uncertainty in their observation likelihood model. Using those prior, the agent can leverage the Active Inference scheme to determine where to direct itself to maximise its objective (e.g. forming a comprehensive map of the environment). While previous models such as  [8, 28] adjust their internal model growth to accommodate new patterns of observations, we extend the concept to predicted, un-visisted, areas and generate new states holding no observation. Those unknown states are therefore highly attractive when seeking information gain and largely improve exploration strategy.

### v2

@@ 1 Introduction
Building on these concepts, we introduce a novel model that dynamically learns environmental structure and expands its cognitive map. Integrating visual information and proprioception (inferred body motion), our model constructs locations and connections within its cognitive map. Starting with uncertainty, the model envisions action outcomes, expanding its map by incorporating hypotheses into its generative model, analogous to Bayesian model reduction [9] that grows its model upon receiving new observation, we extend ours upon predicted beliefs. This process allows our model to efficiently navigate and comprehend environmental structures with minimal steps, using an active inference navigation scheme [17]. Compared to the Clone-Structured Graph (CSCG) [24], our model rapidly learns environmental layouts more efficiently. An exploration run is shown in Figure 1, displaying from left to right the full environment, the extracted observation and exploration path and the resulting internal map of the agent.
@@ 3.1 Inference and spatial abstraction
The classical inference scheme heavily relies on past and current experiences to localise the agent within its environment, using observation alone, the agent would be weak to aliased observations at different locations. By combining observation with the agent’s proprioception the model is much more robust in differentiating ambiguous environments. The internal positioning $p_{0}$ is initialised at the start of exploration in the absence of prior information, and is updated as the agent transitions between rooms (e.i., by passing through a door), thus as long as the agent is confident in its current state. the POMDP factor graph showing the association between poses, states and observations is illustrated in Figure 2
@@
Figure 2: factor graph POMDP of our generative model transitioning from past and present (up to time-step $t$) to future (time-step $t+1$). The pose $p_{t}$ is inferred from the previous pose $p_{t-1}$ and the action from policy $\pi$, while the state $s_{t}$ determined by the corresponding observation $o_{t}$ and influenced by the previous state $s_{t-1}$, pose $p_{t}$ and action $a_{t-1}$. Past actions and observations are assumed observable, indicated by a blue colour. In the future, the actions are defined by a policy $\pi$ influencing the new states and position in orange and new predictions in grey.
@@ 3.3 Incorporating spatial dynamics in model parameters
To extend the internal map (our state space), we propose a novel approach where the agent predicts one-step policy outcomes in all directions considering detected obstacles. $B_{p}$ can expand its position dimension given a motion and $A_{p}$ considers the probability of being at a given state given the position. When we predict a new position given no obstacle in a direction, $B_{p}$ expands. If the expected motion leads to an un-visited location, $s_{t+1}$ does not exist in the model. The state is undefined while the position $p_{t+1}$ is certain, therefore all Markov matrices are expected to grow in their state dimension to match this new prediction. This process enables the dynamic expansion of the dimensions of both the observation and position likelihoods ($A_{o}$, $A_{p}$) and state transition ($B_{s}$) to consider the novel state $s_{t+1}$ in a process equivalent to [9]. Subsequently, the state transition probability ($B_{s}$) and position likelihood ($A_{p}$) can be updated through the same equation 4, here shown with a transition matrix.
$B_{\pi}=Q(s_{t+1}|s_{t},\pi)Q(s_{t})*B_{\pi}*learning\_rate$ (4)
With $s_{t+1}$ being a new state if there are no obstacles detected and a new position is predicted or the same state $s_{t}$ otherwise. The learning rate is set higher for experimented transitions than imagined transitions such that we form new connections weaker toward expected places compared to visited places as we would expect from animal synaptic learning [6]. $A_{o}$ has grown in its state dimension, however, it lacks information regarding the specific observation $o_{t+1}$ expected in that location, resulting in a uniform distribution for that state. Such areas exhibit high uncertainty in their observation likelihood model. Using those prior, the agent can leverage the Active Inference scheme to determine where to direct itself to maximise its objective (e.g. forming a comprehensive map of the environment). While previous models such as  [9, 29] adjust their internal model growth to accommodate new patterns of observations, we extend the concept to predict, un-visited, areas and generate new states holding no observation. Those unknown states are therefore highly attractive when seeking information gain and largely improve exploration strategy.